Consider a subset alpha={x+x2,1+x2,1 2x+2x2}ofP2(R). (a) Show that alpha is a basis for P2(R). (b) For f(x) = 1 + x + x2 2 P2(R), find its coordinator vector [f] alpha with respect to alpha. (c) Let = {1, x, x2} be the standard basis for P2(R), and let f(x) = a + bx + cx2 and g(x) = p+qx+rx2 be the elements of P2(R) and k 2 R. Prove that [f+g] = [f] +[g] and [kf] = k[f] and conclude that the map T : P2(R) ! R3 defined by T(f) = [f] for f 2 P2(R), is linear.Consider a subset x= 4x+2, 1+x? 1-2x + 2x²} of P2(R? (o show that a is a basis for P₂ (R) i) For f(x) = 14x48²€ P2 (R), find
Answer #1

º d = { a+a, ita, 1-2a + 2** } E P(R) let G, C, C3 are scalars in IR site | © (x + 3) + , (+ ) + . ( - 21 + 2) = 0 >(+03) ++ C + 263 1 3-2, + 6 + 2 (1-(2) = 1 3 = 1 > → - e, + 2 - 2 C2 - 3 = - 4 sesa 3 - 262 = 3-3 = 1 the coordinate [f] Wort 2 is H